OpenStudy (anonymous):

Will Fan And Medal Given f(x) = 3x − 10 and g(x) = 12x + 8, solve for (f + g)(x) and select the correct answer below.

OpenStudy (anonymous):

(f + g)(x) = 15x − 2 (f + g)(x) = 9x − 2 (f + g)(x) = 15x + 2 (f + g)(x) = 36x2 − 120x

OpenStudy (anonymous):

plz help